OpenStudy (anonymous):

-3x+2y=10, 4x-7y=4 solve by elimination

OpenStudy (anonymous):

{(−6,−4)}

OpenStudy (anonymous):

I solved it by using matrices by elimination

OpenStudy (anonymous):

x=−6 y=−4

OpenStudy (anonymous):

multiply -3x+2y=10 by 4 multiply 4x-7y=4 by 3 -12x+8y=40 12x-21y=12 combine to get -13y=52 solve for y to get -4 plug -4 into either equation to get x 4x-7(-4)=4 4x+28=4 x=-6